chuckthedonut:

On a map, ½ inch represents 4/5 mile. What distance on the map represents 1 mile? _____ inches on map = 1 mile Show your work.

chuckthedonut:

1 attachment
chuckthedonut:

@shadow can you help?

Swirxy:

o here is how we are going to know the answer. Since 1/2 mile is equal to 4/5 mile, so we will solve this using proportions. 1/2 inches : 4/5 mile : ? Missing: _____ ‎work.

KyledaGreat:

In this case, we can use the ratio method to determine the distance on the map that represents 1 mile. We know that ½ inch represents 4/5 mile, which can be written as: 1/2 inch / 4/5 mile To find out how many inches on the map represent 1 mile, we can set up the following proportion: 1/2 inch / 4/5 mile = x inches / 1 mile where x is the number of inches on the map that represents 1 mile. To solve for x, we can cross-multiply and simplify: 1/2 inch * 1 mile = 4/5 mile * x inches 1/2 mile = 4/5 mile * x inches x inches = (1/2 mile) / (4/5 mile) x inches = (1/2) / (4/5) x inches = 5/8 inch Therefore, what _____ inch on the map represents 1 mile?